OpenStudy (anonymous):

The equation of line CD is y = −2x − 2. Write an equation of a line parallel to line CD in slope-intercept form that contains point (4, 5).

OpenStudy (anonymous):

y = −2x + 13 y = negative 1 over 2x + 7 y = negative 1 over 2x + 3 y = −2x − 3

OpenStudy (alivejeremy):

The equation of line CD is y = −2x − 2. Write an equation of a line parallel to line CD in slope-intercept form that contains point (4, 5). The slope of line parallel to line CD is m = -2. (y - y1) = m(x - x1) y - 5 = -2(x - 4) y - 5 = -2x + 8 y = -2x + 8 + 5 y = -2x + 13

OpenStudy (anonymous):

Thankyou!